OpenStudy (anonymous):

Let f(x) = x/(x-6). Find a function y=g(x) so that (f o g)(x)=2x.

OpenStudy (anonymous):

I've tried plugging in equations. I assume there is a systematic way to do it.

OpenStudy (anonymous):

f^-1 (2x)

OpenStudy (anonymous):

12x/(1-2x)

OpenStudy (amistre64):

\[\frac{x}{x-6} = 2x\] solve for 'x' and make that = g(x)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

maybe work lol

OpenStudy (anonymous):

i dont think amistre

OpenStudy (amistre64):

0 = 2x^2 -13x 0 = x(2x -13) x = 0 or x = 13/2

OpenStudy (amistre64):

its just a shot in the dark ;)

OpenStudy (anonymous):

(fog)(x) = 2x so taking f^-1(x) on both sides g(x) = f^-1(2x)

OpenStudy (anonymous):

so just find f^-1 and plug in 2x in place of x

OpenStudy (amistre64):

yeah, but thats only if you wanna try it the easy way ... thats just not how I roll lol

OpenStudy (anonymous):

him im getting -12x/(1-2x)

OpenStudy (anonymous):

Are you confident on that? I am just confused at this point.

OpenStudy (anonymous):

ill show you the work: x/(x-6)=y x=y(x-6) x=yx-6y x-yx=-6y x(1-y)=-6y x=-6y/(1-y) f^-1(x)=-6x/(1-x) f^1(2x)=-12x/(1-2x)

OpenStudy (anonymous):

you can plug in and see if it works too I geuss

OpenStudy (anonymous):

Oh thanks man. I should have known this. Sub the x's for the y's. Then plug it in. Thanks again.

OpenStudy (anonymous):

no problem, i plugged it in and checked and I think it works